a cylindrical shaped water storage tank with diameter 60 feet and height 20 feet need to be painted on the outside. If the tank is on the ground, find the surface area that needs painting
Answer by stanbon(75887)

a cylindrical shaped water storage tank with diameter 60 feet and height 20 feet need to be painted on the outside. If the tank is on the ground, find the surface area that needs painting
-------------
Area of the side of the tank: A1 = 2(pi)rh = 2(pi)30'*20' = 1200pi sq ft.
----------
Area of the top = (pi)r^2 = (pi)(20)^2 = 400(pi) sq.ft.
-------------
Total Area = 1600(pi)sq ft.
==============
Cheers,
Stan H.